# Particle transport; Dynamic Monte Carlo; Global variance reduction; Weight windows; Hybrid methods; Moment equations

### v1

@@ 5.3 Numerical Solution of HLOSM Equations
In this section, we analyze the hybrid solution of the HLOSM equations used as an auxiliary numerical solution to define weight window centers. The weight windows are updated $u_{ww}^{n}=3$ times with update fractions $f_{p}=\frac{p}{u_{ww}^{n}+1}$ and hence $f_{1}=0.25,f_{2}=0.5,f_{3}=0.75$. The values of $\rho$ and $\epsilon_{min}$ are $1.25$ and $10^{-3}$ respectively, and the number of source particles used is $H^{n}=10^{4}$. Figure 10 presents the hybrid solution of the HLOSM equations, $\phi_{hlosm}$ and the Monte Carlo time-crossing tally solution computed using the HWW algorithm, $\llbracket\phi\rrbracket^{HWW}$ (Eq. (17)). This figure shows hybrid solutions computed from the HLOSM equations without filtering of closures and time step initial conditions ($\phi^{HWW}$), and using filtering techniques: (a) the MA filter with $k=3$ (${\phi}^{HWW-MA}$), (b) the Fourier filter with $\varkappa=30$ (${\phi}^{HWW-F}$). At time step $n=2$, the effects of discretizing the $\delta$-function source are still present, and they are not improved by filtering. At later times $n=4,8$ the unfiltered hybrid solution has similar or less error than the Monte Carlo solution. The filtered hybrid solutions at $n=4,8$ achieve lower relative error than either unfiltered hybrid or Monte Carlo solutions. Figure 11 shows the relative $L_{2}$-norm of errors in ${\phi}_{hlosm}^{n}$ and $\llbracket\phi\rrbracket^{n}$ as functions of time step. By this measure, after $n=2$ the hybrid solution is more accurate than the Monte Carlo solution, and the filtered hybrid solutions are even better. The computationally cheap MA filter performs best overall.
Figure 12 shows the function $F^{n}$ which defines the closure in the HLOSM equations. This figure demonstrates the cell-average values of $F^{n}$ at time steps $n=5,10$ computed by Monte Carlo as well as this closure function after applying filter techniques. HWW denotes the unfiltered closure, HWW-MA the MA filter, and HWW-F the Fourier filter. The parameters of the MA and Fourier filters are $k=3$ and $\varkappa=30$, respectively. At the time step $n=5$, both filters lead to functions with similar large-scale shape. The Fourier filter performs better in removing statistical noise at this stage. At the later times as the particles spread out, the stochastic noise increases. At $n=10$ the filters with the given parameters generate functions with different structures while reducing significantly the noise in the closure.
@@
(c) ${\phi}_{hlosm}^{n}$ and $\llbracket\phi\rrbracket^{n}$ at $n=4$
(d) Relative errors at $n=4$
(e) ${\phi}_{hlosm}^{n}$ and $\llbracket\phi\rrbracket^{n}$ at $n=8$
(f) Relative errors at $n=8$
Figure 10: Numerical solutions $\phi_{hlosm}^{n}$, $\phi_{mc}^{n}$ and their relative errors
@@ 5.4 Numerical Solution of Monte Carlo Algorithm with Hybrid Weight Windows
We now present numerical solutions of Monte Carlo algorithms with (a) HWW using ${\phi}_{hlosm}^{n}$, (b) the lagged weight windows (LWW) defined with the Monte Carlo solution $\langle\phi\rangle^{n-1}$ from the previous time step, and (c) the analog Monte Carlo without weight windows. We also include the versions of the HWW algorithms with MA and Fourier filtering, denoted HWW-MA and HWW-F respectively. The test problem is solved with $H_{p}^{n}=10^{4}$ source particles and the same weight window parameters as Sec. 5.3; $\rho=1.25$, $\epsilon_{min}=10^{-3}$, and $u_{ww}^{n}=3$ with $f_{p}=\frac{p}{u_{ww}^{n}+1}$.
@@
Figure 13 shows three different Monte Carlo solutions $\langle\phi\rangle^{n}$ versus spatial position at time steps $n=4,8,16$. No filtering techniques were used by the HWW algorithm. The effects of weight windows on Monte Carlo particle distribution are illustrated in Figure 14 which presents the number of particle tracks per source particle used by each method on these time steps. The relative standard deviation $\frac{\sigma_{i}^{n}}{\phi_{i}^{n}}$ of the Monte Carlo solutions versus spatial position is presented in Figure 15. At $n=4$, the particle tracks of the LWW algorithm are very non-uniform, leading to a higher stochastic error in the center. The lagged solution under-predicts the wave front position, causing more splitting and increased particle track density in this region. By $n=8$, when solution is changing less, the effect persists but to a lesser degree. The HWW algorithm does not generate as many particles at the wave front as the lagged weight windows, but particles still get much further than those of the analog Monte Carlo. The solution of the HWW algorithm has the most uniform relative standard deviation. The variance of the analog Monte Carlo solution is lowest where the most physical particles are, and the opposite appears true for the solutions of the LWW algorithm. Here the HWW algorithm achieves a balance, that leads to a high quality solution everywhere. The solutions presented in Figure 13 have similar wave front positions for all methods at $n=4,8$. At $n=16$, the solutions are similar for $|x|\leq 10$. Near the wave front, the HWW and LWW solutions are more accurate than the analog solution.
@@
To compute a scalar measure of error on each timestep we use two $L_{2}$-norms. First, we define the traditional $L_{2}$-norm,
$f||_{L_{2}}=\sqrt{\sum_{i=1}^{I}f_{i}^{2}\Delta x_{i}}\,.$ (24)
Second, to evaluate performance in the regions with relatively low particle density, we define a modified $L_{2}$-norm given by
$f||_{L_{2}}^{\phi^{*}}=\sqrt{\sum_{i:\phi_{i}<\phi^{*}}f_{i}^{2}\Delta x_{i}}\,.$ (25)
where $\phi^{*}=10^{-3}$. In Figure 16 the relative $L_{2}$-norms of the error in the Monte Carlo solution, $\langle\phi\rangle^{n}$, are shown versus time. These results indicate that using LWW modifies the particle distribution in a way that increases the relative error in $L_{2}$-norm during the first stage of the problem, and that this error propagates in time. The HWW algorithms achieve similar relative $L_{2}$-norm error to analog Monte Carlo, all less than $13\%$. The relative modified $L_{2}$-norms shown in Figure 17 confirm that the Monte Carlo solution error in low-flux regions decreases when using weight windows. HWW with filtering techniques improve Monte Carlo solution in low-density subdomains at the early stage. All HWW methods improve the solution in the later stage.
[…]
To analyze computational efficacy of the hybrid weight windows we use (a) the figure of merit (FOM) based on the standard deviation, $\sigma^{n}_{i}$, in the $i^{th}$ spatial cell at the $n^{th}$ time step which is given by
$FOM_{\sigma,i}^{n}=\frac{1}{\tau^{n}(\sigma^{n}_{i})^{2}}\,$, (26)
where $\tau^{n}$ is the runtime of the time step $n$, and (b) the FOM based on the error defined by
$FOM_{err,i}^{n}=\frac{1}{\tau^{n}(\phi^{n}_{MC,i}-\phi_{ref,i}^{n})^{2}}\,.$ (27)
The $L_{2}$-norms of these figures of merit, namely, $FOM_{err}||_{L_{2}}$ and $FOM_{\sigma}||_{L_{2}}$ versus time step are plotted in Figures 18 and 19, respectively. The figures present the functions for the HWW algorithm with and without filtering techniques, the LWW and analog algorithms. The parameter of the MA filter is $k=3$. The Fourier filter is applied with $\varkappa=30$. The results show that the algorithms with weight windows have smaller $FOM_{err}||_{L_{2}}$ and $FOM_{\sigma}||_{L_{2}}$ compared to the analog Monte Carlo. The HWW algorithm without filtering has greater $FOM_{err}||_{L_{2}}$ than the LWW one for $n\geq 3$. At the initial stage when the rate of solution change is large, $FOM_{\sigma}||_{L_{2}}$ of the HWW algorithm is greater than this figure of merit for the LWW algorithm. Both kinds of filtering improve the figures of merit of the HWW algorithm indicating better performance compared to the LWW algorithm. The largest impact on $FOM_{err}||_{L_{2}}$ is achieved with the MA filter. $FOM_{\sigma}||_{L_{2}}$ of the HWW algorithm is increased during the initial stage of the problem. We note that hybrid solution $\phi_{losm}$ which is used by the HWW algorithm provides sufficiently accurate approximation of solution over the whole interval of the problem. This advantage is important during stages on which solution magnitude and shape vary significantly over a time step. When solution changes slowly and preserving its spatial shape between time steps, the solution from the previous time step applied in the LWW algorithm is a good approximation for the weight window centers. Figures 20 and 21 show the modified norms of the figures of merit $FOM_{\sigma}||_{L_{2}}^{\phi^{*}}$ and $FOM_{\sigma}||_{L_{2}}^{\phi^{*}}$for $\phi^{*}=10^{-3}$, respectively. Figures 20 and 21 confirm that using the filtered hybrid weight windows leads to a slightly more efficient final solution than either unfiltered hybrid or lagged weight windows.
@@
Figure 18: $FOM_{err}||_{L_{2}}$ versus time for the HWW algorithm with and without filtering techniques, the LWW and analog algorithms
Figure 19: $FOM_{\sigma}||_{L_{2}}$ versus time for the HWW algorithm with and without filtering techniques, the LWW and analog algorithms
@@
Figure 20: $FOM_{err}||_{L_{2}}^{\phi^{*}}$ for $\phi^{*}=10^{-3}$ versus time for the HWW algorithm with and without filtering techniques, the LWW and analog algorithms
@@
Figure 21: $FOM_{\sigma}||_{L_{2}}^{\phi^{*}}$ for $\phi^{*}=10^{-3}$ versus time for the HWW algorithm with and without filtering techniques, the LWW and analog algorithms
6 Discussion
@@
We have developed non-analog Monte Carlo methods for time-dependent transport problems with weight windows for global variance reduction. Weight windows are defined with the solution of a hybrid Monte Carlo / deterministic scheme based on the low-order second-moment equations approximated with the second-order discretization in time and space. One of the methods applies filtering techniques to closures and time-step initial conditions of the low-order equations to reduce effect of stochastic noise on the auxiliary solution.
@@
Filtering the Monte Carlo quantities of the low order problem improved the hybrid solution. Both Fourier and moving average filtering reduced noise and preserved the underlying structure of the closure. The moving average filter proved effective and is computationally cheap.
We have shown that constraining particles to an accurate hybrid auxiliary solution provides a more uniform relative variance than analog Monte Carlo or LWW. The hybrid solution computed with unfiltered quantities had relative L2 error that was smaller than the Monte Carlo solution from the same histories. When the filtering techniques were applied the hybrid solution error decreased further. Using HWW led to a more uniform spatial distribution of particles, and therefore relative stochastic error. In problems where the solution is required at both high and low flux regions, HWW enable efficient solution in the low flux regions. We automate the process of defining a deterministic problem and providing the solution to the Monte Carlo solver. By forming a new hybrid problem at each Monte Carlo census accumulated discretization error is eliminated. This also allows for integration with multiphysics codes, where material properties and sources are evaluated at the census and not known beforehand.
This method is limited by the additional cost of the hybrid problem. A large portion of this cost is the tallies needed to define the low order problem. These tallies could be made cheaper or the hybrid problem could be formulated to require less additional tallies. Monotonization techniques can be applied to modify the second-order scheme for discretization of the LOSM equations [17, 18] The developed algorithms can be extended to multidimensional problems.

### v2

@@ 5.4 Numerical Solution of HLOSM Equations
In this section, we analyze the hybrid solution of the HLOSM equations used as an auxiliary numerical solution to define weight window centers. The weight windows are updated $u_{ww}^{n}=3$ times with update fractions $f_{p}=\frac{p}{u_{ww}^{n}+1}$ and hence $f_{1}=0.25,f_{2}=0.5,f_{3}=0.75$. The values of $\rho$ and $\epsilon_{min}$ are $1.25$ and $10^{-3}$ respectively, and the number of source particles used is $H^{n}=10^{4}$. Figures 10, 11, and 12 show the hybrid solution computed using the HWW algorithm, ${\phi}^{HWW}$, the Monte Carlo time-crossing tally solution, $\llbracket\phi\rrbracket^{HWW}$ (Eq. (18)), and their relative errors for the time steps $n=2,4,8$. These figures also include two hybrid solutions computed using filtered initial conditions and closures. One hybrid solution, ${\phi}^{HWW-MA}$, uses the MA filter with $k=3$ and the other, ${\phi}^{HWW-F}$, the Fourier filter with $\varkappa=30$. At time step $n=2$, the effects of discretizing the delta function source are still present, and they are not improved by filtering. At later times $n=4,8$, the unfiltered hybrid solution has similar or less error than the Monte Carlo solution. The filtered hybrid solutions at $n=4,8$ achieve lower relative error than either unfiltered hybrid or Monte Carlo solutions. Figure 13 shows the relative $L_{2}$-norm of errors in ${\phi}_{hlosm}^{n}$ and $\llbracket\phi\rrbracket^{n}$ as functions of time step. By this measure, after $n=2$ the hybrid solution is more accurate than the Monte Carlo solution, and the filtered hybrid solutions are even better. The computationally cheap MA filter performs best overall.
Figure 14 shows the function $F^{n}$ which defines the closure in the HLOSM equations. This figure demonstrates the cell-average values of $F^{n}$ at time steps $n=5,10$ computed by Monte Carlo as well as this closure function after applying filter techniques. HWW denotes the unfiltered closure, HWW-MA the MA filter, and HWW-F the Fourier filter. The parameters of the MA and Fourier filters are $k=3$ and $\varkappa=30$, respectively. At the time step $n=5$, both filters lead to functions with a similar large-scale shape. The Fourier filter performs better in removing statistical noise at this stage. At later times as the particles spread out, the stochastic noise increases. At $n=10$ the filters with the given parameters generate functions with different structures while reducing significantly the noise in the closure.
@@
Figure 10: Numerical solutions $\phi_{hlosm}^{n}$, $\phi_{mc}^{n}$ and their relative errors at $n=2$
(a) ${\phi}_{hlosm}^{n}$ and $\llbracket\phi\rrbracket^{n}$ at $n=4$
(b) Relative errors at $n=4$
Figure 11: Numerical solutions $\phi_{hlosm}^{n}$, $\phi_{mc}^{n}$ and their relative errors at $n=4$
(a) ${\phi}_{hlosm}^{n}$ and $\llbracket\phi\rrbracket^{n}$ at $n=8$
(b) Relative errors at $n=8$
Figure 12: Numerical solutions $\phi_{hlosm}^{n}$, $\phi_{mc}^{n}$ and their relative errors at $n=8$
@@ 5.5 Numerical Solution of Monte Carlo Algorithm with Hybrid Weight Windows
We now present numerical solutions of Monte Carlo algorithms with (a) the hybrid weight windows (HWW) using ${\phi}_{hlosm}^{n}$, (b) the lagged weight windows (LWW) defined with the Monte Carlo solution $\langle\phi\rangle^{n-1}$ from the previous time step, and (c) the analog Monte Carlo without weight windows. We also include the versions of the HWW algorithms with MA and Fourier filtering, denoted HWW-MA and HWW-F respectively. The parameter of the MA filter is $k=3$. The Fourier filter is applied with $\varkappa=30$. The test problem is solved with $H_{p}^{n}=10^{4}$ source particles and the same weight window parameters as in Section 5.4; $\rho=1.25$, $\epsilon_{min}=10^{-3}$, and $u_{ww}^{n}=3$ with $f_{p}=\frac{p}{u_{ww}^{n}+1}$.
@@
Figure 15 shows three different Monte Carlo solutions $\langle\phi\rangle^{n}$ versus spatial position at time steps $n=4,8,16$. No filtering techniques were used by the HWW algorithm. The effects of weight windows on Monte Carlo particle distribution are illustrated in Figure 16, which presents the number of particle tracks per source particle used by each method on these time steps. The relative standard deviation $\frac{\sigma_{i}^{n}}{\phi_{i}^{n}}$ of the Monte Carlo solutions versus spatial position is presented in Figure 18. At $n=4$, the particle tracks of the LWW algorithm are very non-uniform, leading to a higher stochastic error in the center. The lagged solution under-predicts the wave front position, causing more splitting and increased particle track density in this region. By $n=8$, when the solution is changing less, the effect persists but to a lesser degree. The HWW algorithm does not generate as many particles at the wave front as the lagged weight windows, but particles still get much further than those of the analog Monte Carlo. The solution of the HWW algorithm has the most uniform relative standard deviation. The variance of the analog Monte Carlo solution is lowest where the most physical particles are, and the opposite appears true for the solutions of the LWW algorithm. Here the HWW algorithm achieves a balance, that leads to a high quality solution everywhere. The solutions presented in Figure 15 have similar wave front positions for all methods at $n=4,8$. At $n=16$, the solutions are similar for $|x|\leq 10$. Near the wave front, the HWW and LWW solutions are more accurate than the analog solution.
@@
Figure 18 shows the wavefront position for each of the methods studied as a function of time. This is defined as the outer edge of furthest cell with particles in it. This figure shows that analog Monte Carlo is unable to accurately track the wavefront beyond $5s$, and WW are essential to achieve an accurate wavefront position at later times. The point where the WW algorithms depart from the reference is determined by the $\varepsilon_{min}$ parameter.
@@
Figure 18: Furthest cell with particles versus time for the HWW algorithm with and without filtering techniques, the LWW and analog algorithms
To quantify the accuracy of Monte Carlo solutions, we use the relative standard deviation and the relative error. The average of cell-wise relative standard deviation is defined as
$\overline{RSD}^{n}=\frac{1}{I}\sum_{i=1}^{I}\frac{\sigma_{i}^{n}}{\phi_{i,mc}^{n}}.$ (25)
This measure does not account for regions where the reference solution is nonzero and Monte Carlo particles have not reached. The average cell-wise relative error, defined as,
$\overline{RE}^{n}=\frac{1}{I}\sum_{i=1}^{I}\frac{|\phi_{i,mc}^{n}-\phi_{i,ref}^{n}|}{\phi_{i,ref}^{n}}$ (26)
is an explicit measure of the solution quality. A spatial average of both measures is computed on each timestep and plotted in Figures 19 and 20. In Figure 19 the positive effects of weight windows are seen between $n=4$ and $n=11$, where the more uniform relative standard deviation improves the average. Figure 20 shows that the error caused by particles not making it to the edges of the problem is significant and dominates the analog solution. Also in Figure 20 we can see the LWW solution has increased error caused by the uneven redistribution of particles during early timesteps. The HWW achieve the lowest relative error after $n=4$, because of the more uniform particle distribution and more particles reaching the wavefront.
@@
Figure 19: Average relative standard deviation for the HWW algorithms with and without filtering techniques, the LWW and analog algorithms
@@
Figure 20: Average relative error for the HWW algorithms with and without filtering techniques, the LWW and analog algorithms
To measure the efficiency of each algorithm, we define two figures of merit (FOM). One FOM is computed using relative standard deviation and one using relative error defined, respectively, as follows:
$FOM_{\sigma}^{n}=\frac{1}{\tau^{n}\left[\frac{1}{I}\sum_{i=1}^{I}\left(\frac{\sigma_{i}^{n}}{\phi_{i,mc}^{n}}\right)^{2}\right]}\,$, (27)
$FOM_{Error}^{n}=\frac{1}{\tau^{n}\left[\frac{1}{I}\sum_{i=1}^{I}\left(\frac{|\phi_{i,mc}^{n}-\phi_{i,ref}^{n}|}{\phi_{i,ref}^{n}}\right)^{2}\right]}\,$, (28)
where $\tau^{n}$ is the runtime of the $n^{th}$ census.
In Figures 22 and 22 are the FOMs for the HWW algorithms with and without filtering, LWW and analog Monte Carlo. The LWW and analog methods have shorter runtimes, since they require less tallies. This causes their $FOM_{\sigma}$ to be greater than the hybrid methods. In Figures 24 and 24 the FOMs for HWW and LWW are normalized by the analog FOM. Figure 24 shows that weight window methods generally improve $FOM_{Error}$ and the filtered hybrid methods improve $FOM_{Error}$ between $n=2$ and $n=15$. The decrease in $FOM_{Error}$ in the early time steps comes from splitting that results in additional particles being simulated, increasing runtime. During this period of time, the analog particle distribution is sufficient to capture the solution accurately, so weight windows decrease efficiency. The HWW algorithm with moving average filter achieves the greatest $FOM_{Error}$ increase of the hybrid methods, with an average of 1.25x improvement over analog. HWW with Fourier filtering, no filtering, and LWW achieved averages of 1.15x, 0.92x, 1.30x respectively.
Figure 21: $FOM_{\sigma}$ for the HWW algorithm with and without filtering techniques, the LWW and analog algorithms
Figure 22: $FOM_{Error}$ for the HWW algorithm with and without filtering techniques, the LWW and analog algorithms
Figure 23: Normalized $FOM_{\sigma}$ for the HWW algorithm with and without filtering techniques and the LWW algorithm
Figure 24: Normalized $FOM_{Error}$ for the HWW algorithm with and without filtering techniques, the LWW algorithm
6 Conclusion
@@
We have developed non-analog Monte Carlo methods for time-dependent transport problems with weight windows for global variance reduction. Weight windows are defined with the solution of a hybrid Monte Carlo / deterministic scheme based on the low-order second-moment equations approximated with the second-order discretization in time and space. One of the methods applies filtering techniques to closures and time-step initial conditions of the low-order equations to reduce the effect of stochastic noise on the auxiliary solution.
@@
Filtering the Monte Carlo quantities of the low order problem improved the hybrid solution. Both Fourier and moving average filtering reduced noise and preserved the underlying structure of the closure. The moving average filter proved be effective and is computationally cheap.
We have shown that constraining particles to an accurate hybrid auxiliary solution (HWW) provides a more uniform relative variance than analog Monte Carlo or LWW. The hybrid solution computed with unfiltered quantities had relative L2 error that was smaller than the Monte Carlo solution from the same histories. When the filtering techniques were applied, the hybrid solution error decreased further. Using HWW led to a more uniform spatial distribution of particles, and therefore relative stochastic error. In problems where the solution is required at both high and low flux regions, HWW enable efficient solution in the low flux regions. The relative error based figure of merit for HWW showed improvement over analog Monte Carlo. We automate the process of defining a deterministic problem and providing the solution to the Monte Carlo solver. By forming a new hybrid problem at each Monte Carlo census, accumulated discretization error is eliminated. This also allows for integration with multiphysics codes, where material properties and sources are evaluated at the census and not known beforehand.
This method is limited by the additional cost of the hybrid problem. A large portion of this cost is the tallies needed to define the low order problem. These tallies could be made cheaper or the hybrid problem could be formulated to require less additional tallies. Monotonization techniques can be applied to modify the second-order scheme for discretization of the LOSM equations [23, 24] The developed algorithms can be extended to multidimensional problems to further realize the benefits of the self-adjoint second moment operator and filtering.